A cough syrup (density 1.25 g/mL) is labelled as containing 16% (v/v) ethanol (density 0.789 g/mL). Express the concentration of the ethanol as percentage (w/w) and as mol/L.              

1
Expert's answer
2021-01-18T01:31:46-0500

d = m/v => w/w = (1.25 g/mL/ 0.789 g/mL )%

= 1.584%

Molar mass of ethanol = 46.07 g/mol

Molar concentration(mol/dm3) = density/molar mass

= 0.789 g/mL / 46.07 g/mol

= 0.017 * 10-3mol/L